# MambaPEFT: Exploring Parameter-Efficient Fine-Tuning for Mamba

### v1

@@ B.1 Language Tasks
As to the learning rate, we use suitable values for each method. It is set to 5e-5 for the full fine-tuning of Mamba and Pythia (Biderman et al., 2023). For SLL LoRA on Mamba, we follow the original learning rate per model size (Halloran et al., 2024). We also use a roughly tuned learning rate for our PEFT methods obtained through the ablation studies in Section A.4. The configurations are shown in Table 5.
@@
Method Mamba 160M Mamba 370M Mamba 790M Mamba 1.4B
@@
LoRA on Pythia 1e-5 1e-4 1e-5 1e-5
For the hyperparameters of each PEFT method, we use the optimal values obtained through the ablation studies on the VTAB-1K dataset. For Pythia, We apply LoRA with $r=8$ to dense, query_key_value, dense_4h_to_h, and dense_h_to_4h layers.
@@ B.2 Hybrid PEFT Search
Since optimizing over all 19 tasks is computationally expensive, we optimize over the average accuracy of 5 tasks in the representative subset of VTAB-1k (Zhai et al., 2019): Caltech101, EuroSAT, Pets, Camelyon, and Resisc45. By processing five tasks in parallel on one A100 GPU, one trial can be completed in 20 to 30 minutes, with minimal dependency on the type and size of the applied PEFT methods.
As mentioned in the Section 3.4, the first step of our search verifies the effectiveness of each PEFT method and seeks the preferable combination with minimum trainable parameters. Specifically, we search for 100 trials using TPE algorithm (Bergstra et al., 2011) about whether to use each PEFT method with the fixed hyperparameters shown in step 1 of Table 6. As a result, it turns out that the combination of eight PEFT methods, CLS-token-tuning, A-tuning, Affix-tuning, LoRA(out_proj), LoRA(in_proj), LoRA(dt_proj), LoRAp(X), and LoRAp(B), is the best.
In the second step, the hyperparameters for the eight methods are searched. In addition, a search dimension is added which selects a PEFT method that should be removed or decides not to remove any at all. This is intended to eliminate less important PEFT methods and allocate more parameters to the more significant ones. The specific search space is shown in step 2 of Table 6.

### v2

@@ A.4 Ablation studies on language tasks
[…]
[mode=buildnew]fig/sun
[…]
[mode=buildnew]fig/reti
[…]
[mode=buildnew]fig/dmlab
Figure 7: Task-specific relationship between size of training data and performance of PEFT in Vim. (a) is on Sun397, (b) is on Retinopathy, and (c) is on DMLab task.
Base Model Method #Params (K) Natural Specialized Structured Average
VMamba-S Full 49,379 66.4 81.8 62.6 68.0
Linear 18 67.9 73.6 34.2 54.9
Conv2d-tuning 160 75.8 83.3 58.9 70.3
Affix-tuning(w/o proj) 198 77.3 82.7 59.7 71.0
Additional-scan 395 76.4 83.3 60.2 71.0
ParallelAdapter 774 77.8 84.5 63.7 73.2
LoRA(out_proj) 1,529 77.9 85.9 65.6 74.4
LoRAp(X) 1,529 78.2 84.5 65.2 74.0
VMamba-B Full 87,533 66.8 81.4 60.1 67.1
Linear 24 68.4 79.1 30.2 54.6
Conv2d-tuning 189 78.0 84.1 63.4 73.1
Affix-tuning(w/o proj) 276 77.5 83.3 59.2 71.0
Additional-scan 527 77.0 83.8 60.1 71.3
ParallelAdapter 1,032 78.6 85.1 63.3 73.5
LoRA(out_proj) 2,039 78.6 85.7 64.9 74.3
LoRAp(X) 2,039 78.3 85.3 65.2 74.3
Table 5: Comparison of methods for VMamba. In VMamba, LoRAp(X) is the same as LoRA(in_proj).
A.5 Effect of data size
Table 1 demonstrates that Partial LoRA performs better in image tasks. On the other hand, Table 3 shows that Affix-tuning and Additional-scan achieve competitive accuracy with Partial LoRA despite fewer parameters in language tasks. The major differences between the two experiments are the modality and the amount of training data. In the image experiment, fine-tuning was performed with 1K images, while in the language experiment, fine-tuning was conducted with 170K data. To investigate whether the differences in trends are due to the amount of data, we conducted experiments to scale the amount of training data in the image tasks. In detail, we used official test data of VTAB-1K as training data and vice versa. We chose Sun397, Retinopathy, and DMLab from each category of VTAB-1K, which contains a large number of images for this experiment. fig. 7 shows that, as the data size increases, the best methods become Additional-scan for (a), Affix-tuning for (b), and LoRAp(X) for (c). These experiments suggest that LoRAp(X) excels with smaller data sizes, but with larger data sizes, the optimal method depends on the specific task. We suggest selecting a method that balances the trade-off between performance and parameters based on the task or searching HybridPEFT.
A.6 Effect of model size
Section A.3 contains experimental results for Vim and ViT with size variations. From Vim-Tiny to Vim-S, the overall accuracy is improved. However, from Vim-S to Vim-B, there is little improvement, whereas ViT-B improved from ViT-S. Because this phenomenon is not observed in vanilla Mamba for language tasks (see table 3), it seems it is not due to PEFT methods for Mamba. This suggests that large Vim/Mamba models might face potential challenges when fine-tuning with limited data. Note that even with Vim-B, PEFT still significantly outperforms full fine-tuning and achieves equivalent or higher accuracy than state-of-the-art PEFT with ViT-B, indicating that PEFT for Vim remains effective.
A.7 Experiments with VMamba
To investigate the generality of PEFT in Mamba for visual tasks, we conduct experiments using VMamba (Liu et al., 2024b), a variant other than Vim. VMamba made several improvements over the original Mamba. It eliminated the state dimension (i.e., reduced the state dimension to 1), introduced a hierarchical structure similar to CNNs, replaced causal 1D convolutions with 2D convolutions, removed the Gated MLP structure, and substituted half of the Mamba blocks with MLP blocks. Due to these changes, there are some differences in trends, but PEFT also works well with VMamba, as shown in Table 5. The accuracy is higher than that of Vim, as the original VMamba performs better on ImageNet1K evaluation. Similar to the results with Vim, partial LoRA shows superior performance. Additionl-scan and Partial-tuning show a superior trade-off between trainable parameters and performance. Furthermore, there was no significant performance improvement from VMamba-S to VMamba-B, indicating saturation, which is also consistent with Vim. These experimental results suggest that the phenomenon of performance saturation with PEFT in Mamba for visual tasks, once the model size reaches a certain threshold, is general regardless of variation.
A.8 Direction of the Future Work
Our goal is to explore PEFT for Mamba: specifically, how to tune and where to tune. By using a simple architecture PEFT methods, we aim to minimize extraneous influences, allowing evaluation results to be universally informative for various future tuning strategies. The proposed Affix-tuning and Additional-scan are also designed as simple as possible, although we believe they possess significant novelty and usefulness. Based on our findings, future work should be able to focus on effective PEFT research without searching a lot. To verify this, we conducted an experiment by replacing LoRA with the latest DoRA (Liu et al., 2024a). As shown in Table 6, we were able to achieve even higher accuracy than LoRAp(X), which already surpasses the accuracy of state-of-the-art dedicated PEFT for Transformers. We hope that our exploration will contribute to the future evolution of PEFT for Mamba.
Table 6: Replacing simple structure PEFT methods we intentionally used with dedicated PEFT methods.
VTAB-1K with Vim-S Commonsense reasoning with Mamba 1.4B
Method Natu. Spec. Struc. Avg. BoolQ PIQA SIQA HellaS WinoG ARC-e ARC-c OBQA Avg.
LoRAp(X) 76.6 83.9 60.8 71.5 63.1 73.5 42.7 57.7 61.6 60.4 32.9 37.4 53.7
DoRAp(X) 76.9 84.1 60.6 71.6 63.9 73.8 43.2 57.9 62.1 60.4 32.4 37.4 53.9
@@ B.1 Language Tasks
As to the learning rate, we use suitable values for each method. It is set to 5e-5 for the full fine-tuning of Mamba. For SLL LoRA on Mamba, we follow the original learning rate per model size (Halloran et al., 2024). We also use a roughly tuned learning rate for our PEFT methods obtained through the ablation studies in Section A.4. The configurations are shown in Table 7.
@@
Method Mamba 130M Mamba 370M Mamba 790M Mamba 1.4B
@@
Method Pythia 160M Pythia 410M Pythia 1B Pythia 1.4B
Full 5e-5 5e-5 5e-5 5e-6
LoRA 1e-5 1e-4 1e-5 1e-5
For the hyperparameters of each PEFT method, we use the optimal values obtained through the ablation studies on the VTAB-1K dataset. For Pythia (Biderman et al., 2023), We apply LoRA with $r=8$ to dense, query_key_value, dense_4h_to_h, and dense_h_to_4h layers.
@@ B.2 Hybrid PEFT Search
In the first step, we explore only whether to use each PEFT method. As mentioned in the Section 3.4, the first step of our search verifies the effectiveness of each PEFT method and seeks the preferable combination with minimum trainable parameters. We rely on the following hypothesis; effective PEFT methods work positively even when the number of trainable parameters is small, while when the number of parameters increases beyond a certain level, the pre-trained model collapses and is negatively affected. As table 8 shows, simply combining PEFT or increasing the rank of LoRA decreases performance. Therefore, the hyperparameters of each PEFT method are set as its trainable parameters become minimum. In addition, the proposed regularization for Partial-tuning PEFT methods is applied so as not to corrupt the pre-trained model. In this step, a combination of PEFT methods that work positively is selected greedy while keeping the total degrees of freedom of the PEFT methods within the degrees of freedom that do not destroy the pre-trained model. We search for 100 trials using TPE algorithm (Bergstra et al., 2011) about whether to use each PEFT method with the fixed hyperparameters shown in step 1 of Table 9. As a result, it turns out that the combination of eight PEFT methods, CLS-token-tuning, A-tuning, Affix-tuning, LoRA(out_proj), LoRA(in_proj), LoRA(dt_proj), LoRAp(X), and LoRAp(B), is the best.
In the second step, the hyperparameters for the eight methods are searched. In addition, a search dimension is added, which selects a PEFT method that should be removed or decides not to remove any at all. This is intended to eliminate less important PEFT methods and allocate more parameters to the more significant ones. The specific search space is shown in step 2 of Table 9.
This two-step approach reduces the search space to only $2^{N_{\mathrm{all}}}$ for the first step and $(N_{i}+1)\prod_{p\in P_{i}}d_{p}$ for the subsequent steps, compared to a simultaneous search space with $2^{N_{\mathrm{all}}}\prod_{p\in P_{\mathrm{all}}}d_{p}$. $N_{\mathrm{all}}$ is the number of PEFT methods, and $N_{i}(\leq N_{\mathrm{all}})$ is the number of methods chosen in the previous search step. The $P_{\mathrm{all}}$ is the set of all the hyperparameters and $P_{i}(\subseteq P_{\mathrm{all}})$ is the subset used in the chosen methods. Also, we represent the search space size along a parameter $p$ as $d_{p}$.
Since optimizing over all 19 tasks is computationally expensive, we optimize over the average accuracy of 5 tasks in the representative subset of VTAB-1k (Zhai et al., 2019): Caltech101, EuroSAT, Pets, Camelyon, and Resisc45. By processing five tasks in parallel on one A100 GPU, one trial can be completed in around 20 minutes, with minimal dependency on the type and size of the applied PEFT methods.
@@
Method Rank Avg.
LoRA(in_proj + out_proj) 8 71.33
LoRA(in_proj + out_proj) 16 70.68
LoRA(in_proj + out_proj) 32 69.77
LoRA(ALL) 8 71.02
LoRA(ALL) 16 70.42
LoRA(ALL) 32 68.74
LoRAp(X) 64 71.52
Table 8: Average performance of VTAB-1K when combining multiple LoRA.